# Plana3R: Zero-shot Metric Planar 3D Reconstruction via Feed-Forward Planar Splatting

### v1

@@ 1 Introduction
As indoor scenes are typically rich in planar structures such as floors, ceilings, and walls, as well as planar furniture like tables and cabinets, planar primitives are well-suited representations for the accurate 3D reconstruction of indoor scenes. As a result, there has been significant interest among the research community in planar 3D reconstruction in recent years. Planar reconstruction approaches include feedforward solutions in monocular [39, 16, 26, 23] and two-view [11, 1, 27] settings, and per-scene optimization approaches [28, 37, 3, 9] that leverage posed multi-view inputs with the assistance of the feedforward models were studied. However, these approaches face two key limitations:
@@
  • Pose dependence for optimization-based methods: Per-scene optimization techniques [28, 9, 37, 3] depend on accurately posed multi-view, densely-captured images, which are not always available and would lead to either undermining reconstruction quality or limited usage scenarios. Furthermore, those methods cannot handle several sparse pose-free views or just a pair of pose-free views.
@@ 3.3 Training Losses and Training Strategies
where $\ast\in\{\text{low},\text{high},\text{selected},\}$ and $\beta_{1},\beta_{2}$ balance the loss magnitudes for stable training.
@@ 4.2 Datasets
Since Plana3R targets structured indoor scenes, we train it on a combination of four public indoor-scene datasets: ScanNetV2 [4], ScanNet++ [38], ARKitScenes [5], and Habitat [22]. From these datasets, we construct approximately four million image pairs. Pseudo GT normal maps are generated using Metric3Dv2 [10], while GT depth maps are directly taken from the datasets. For evaluation, we use ScanNetV2, Matterport3D [2], NYUv2-Plane [19, 15], and Replica [25] as test sets. Among these test sets, except for ScanNetV2, the remaining three datasets demonstrate the generalization capability of our model across different datasets. We also evaluate on 7Scenes [24] in Sec.˜A.1.
@@ 4.3.2 Metric Monocular Depth Estimation
Directly rendering depth from the predicted planar primitives provides an effective means of evaluating the quality of our geometric fitting. This process does not require merging the primitives and can be performed with a single feed-forward pass. We evaluate the metric monocular depth estimation using the NYUv2-Plane dataset. NYUv2-Plane provided by [15] is a planar variant of the original NYUv2 dataset [19]. For our experiments, we use the image splits defined in [23, 15] with 654 test frames. We input two identical images into Plana3R and use the CUDA-based differentiable planar primitive rendering module in PlanarSplatting to generate a depth map of the predicted primitives from one of the views.
@@
Table 2: Quantitative comparison of metric monocular depth estimation on the NYUv2-Plane [19, 15] dataset. The best results are in bold.
@@
$\delta_{1}$ $\uparrow$ 53.0 61.3 71.8 63.3 74.2 80.6 86.4
$\delta_{2}$ $\uparrow$ 78.3 87.2 93.1 88.2 94.2 94.7 96.3
$\delta_{3}$ $\uparrow$ 90.4 95.8 98.3 96.1 99.0 98.5 98.4
@@
Single-view 3D Plane Segmentation
Following PlaneRCNN [16], we generate 3D plane GT labels on the Replica dataset [25] by first fitting planes to the GT mesh using RANSAC [8], and then splitting co-planar regions with different semantics into separate plane instances. A total of 498 images are sampled for single-view plane-level segmentation evaluation.
Similar to monocular depth estimation, we input two identical images into Plana3R, then we merge the output primitives. Tab.˜3 and Fig.˜4 show that our Plana3R performs better than the method trained with plane annotation in both plane segmentation and 3D reconstruction. We present more visualization results in the supplementary materials and conduct tests on the 7-Scenes [24] dataset.
Figure 4: Comparisons of single-view plane segmentaion and 3D reconstruction on the Replica [25].
@@ 4.5 Alation Study
ScanNetV2 Reconstruction NYUv2-Plane Depth Estimation
@@ A.1 Extra Results
We present more visualization results here.
@@
To further demonstrate the zero-shot generalization capability of our method in out-of-domain scenes, we evaluate single-view reconstruction and planar segmentation on the 7-Scenes dataset [24]. Although 7-Scenes is a widely used indoor dataset and is very suitable for out-of-domain evaluation, it does not provide official plane segmentation masks. We attempted to generate them via mesh synthesis but found it challenging to obtain high-quality ground truth labels. Therefore, we present the visualization results. As shown in Fig.˜7, our method achieves much better planar segmentation and reconstruction performance than PlaneRecTR.
@@ A.2 Implementation Details
Performance Analysis on Pair’s Overlap
@@
Our Plana3R predicts a set of 3D planar primitives from the inputs. To further achieve the instances of 3D planes, we merge the predicted primitives according to their spatial connectivity and geometric similarity. In detail, we greedily merge adjacent planar primitives that meet the merging thresholds (projection distance $\leq$ 0.1m and normal angle difference $\leq$ 25 degrees) to obtain the final 3D plane instances.
@@ A.3 Runtime Analysis
A.4 Planar Primtive and Primitive Selection

### v2

@@ 1 Introduction
As indoor scenes are typically rich in planar structures such as floors, ceilings, and walls, as well as planar furniture like tables and cabinets, planar primitives are well-suited representations for the accurate 3D reconstruction of indoor scenes. As a result, there has been significant interest among the research community in planar 3D reconstruction in recent years. Planar reconstruction approaches include feedforward solutions in monocular [40, 16, 27, 24, 18, 42] and two-view [11, 1, 28] settings, and per-scene optimization approaches [29, 38, 3, 9] that leverage posed multi-view inputs with the assistance of the feedforward models were studied. However, these approaches face two key limitations:
@@
  • Pose dependence for optimization-based methods: Per-scene optimization techniques [29, 9, 38, 3] depend on accurately posed multi-view, densely-captured images, which are not always available and would lead to either undermining reconstruction quality or limited usage scenarios. Furthermore, these methods cannot handle several sparse pose-free views or just a pair of pose-free views.
@@ 3.3 Training Losses and Training Strategies
where $\ast\in\{\text{low},\text{high},\text{selected}\}$ and $\beta_{1},\beta_{2}$ balance the loss magnitudes for stable training.
@@ 4.2 Datasets
Since Plana3R targets structured indoor scenes, we train it on a combination of four public indoor-scene datasets: ScanNetV2 [4], ScanNet++ [39], ARKitScenes [5], and Habitat [23]. From these datasets, we construct approximately four million image pairs. Pseudo GT normal maps are generated using Metric3Dv2 [10], while GT depth maps are directly taken from the datasets. For evaluation, we use ScanNetV2, Matterport3D [2], NYUv2 [20], and Replica [26] as test sets. Among these test sets, except for ScanNetV2, the remaining three datasets demonstrate the generalization capability of our model across different datasets. We also evaluate on 7Scenes [25] in Sec.˜A.1.
@@ 4.3.2 Metric Monocular Depth Estimation
Directly rendering depth from the predicted planar primitives provides an effective means of evaluating the quality of our geometric fitting. This process does not require merging the primitives and can be performed with a single feed-forward pass. We evaluate the metric monocular depth estimation using the NYUv2 [20] dataset. For our experiments, we use the image splits defined in [24, 15] with 654 test frames. We input two identical images into Plana3R and use the CUDA-based differentiable planar primitive rendering module in PlanarSplatting to generate a depth map of the predicted primitives from one of the views.
@@
Table 2: Quantitative comparison of metric monocular depth estimation on the NYUv2 dataset. The best results are in bold.
@@
$\delta_{1}$ $\uparrow$ 53.0 61.3 71.8 63.3 74.2 83.0 86.4
$\delta_{2}$ $\uparrow$ 78.3 87.2 93.1 88.2 94.2 95.6 96.3
$\delta_{3}$ $\uparrow$ 90.4 95.8 98.3 96.1 99.0 98.7 98.4
@@
Single-view 3D Plane Segmentation.
Following PlaneRCNN [16], we generate 3D plane GT labels on the Replica dataset [26] by first fitting planes to the GT mesh using RANSAC [8], and then splitting co-planar regions with different semantics into separate plane instances. A total of 498 images are sampled for single-view plane-level segmentation evaluation. Similar to monocular depth estimation, we input two identical images into Plana3R, and then we merge the output primitives. Tab.˜3 and Fig.˜4 show that our Plana3R performs better than the method trained with plane annotation in both plane segmentation and 3D reconstruction. We present more visualization results in the supplementary materials and conduct tests on the 7-Scenes [25] dataset.
Figure 4: Comparisons of single-view plane segmentation and 3D reconstruction on the Replica [26].
@@ 4.5 Alation Study
ScanNetV2 Reconstruction NYUv2 Depth Estimation
@@ A.1 Extra Results
We present additional visualization results here.
@@
To further demonstrate the zero-shot generalization capability of our method in out-of-domain scenes, we evaluate single-view reconstruction and planar segmentation on the 7-Scenes dataset [25]. Although the 7-Scenes dataset is a widely used indoor dataset and is very suitable for out-of-domain evaluation, it does not provide official plane segmentation masks. We attempted to generate them via mesh synthesis but found it challenging to obtain high-quality ground truth labels. Therefore, we present the visualization results. As shown in Fig.˜7, our method achieves much better planar segmentation and reconstruction performance than PlaneRecTR.
@@ A.2 Implementation Details
Performance Analysis on Pair’s Overlap.
@@
Our Plana3R predicts a set of 3D planar primitives from the inputs. To further achieve instances of 3D planes, we merge the predicted primitives according to their spatial connectivity and geometric similarity. In detail, we greedily merge adjacent planar primitives that meet the merging thresholds (projection distance $\leq$ 0.1m and normal angle difference $\leq$ 25 degrees) to obtain the final 3D plane instances.
@@ A.3 Runtime Analysis
A.4 Planar Primitive and Primitive Selection